# Dynamic-SUPERB Phase-2: A Collaboratively Expanding Benchmark for Measuring the Capabilities of Spoken Language Models with 180 Tasks

### v1

@@ Appendix D HEAR & MARBLE Evaluation Results
| Task | Metric | Whisper- | LTU-AS | SALMONN | SALMONN | Qwen- | Qwen2- | WavLLM | MU- | GAMA |
| LLaMA | 7B | 13B | Audio | Audio | LLaMA | |||||
| Beehive States | LLM-C | 43.00% | 35.00% | 54.00% | 42.50% | 19.00% | 42.50% | 37.00% | 37.50% | 19.00% |
| CREMA-D | LLM-C | 6.70% | 9.30% | 18.80% | 18.10% | 62.50% | 61.10% | 25.60% | 12.40% | 13.20% |
| ESC-50 | LLM-C | 4.80% | 25.30% | 61.00% | 73.40% | 68.80% | 43.90% | 9.80% | 0.40% | 35.30% |
| VoxLingua107 Top10 | LLM-C | 92.18% | 1.95% | 6.28% | 11.01% | 18.00% | 36.11% | 1.85% | 0.00% | 0.10% |
| GTZAN Genre | LLM-C | 7.30% | 9.70% | 37.50% | 42.50% | 29.00% | 29.40% | 9.50% | 26.90% | 7.00% |
| GTZAN Music Speech | LLM-C | 63.28% | 34.38% | 74.22% | 75.78% | 96.88% | 75.00% | 60.94% | 53.13% | 82.81% |
| MAESTRO | NAR | 92.43% | 100.00% | 99.46% | 100.00% | 96.76% | 96.76% | 96.76% | 100.00% | 100.00% |
@@
| Beijing Opera Percussion | LLM-C | 3.39% | 4.24% | 10.17% | 20.76% | 16.10% | 27.12% | 24.15% | 22.03% | 11.44% |
| Mridangam Stroke | LLM-C | 0.10% | 3.80% | 7.60% | 4.90% | 11.50% | 10.80% | 13.70% | 10.30% | 0.10% |
| Mridangam Tonic | LLM-C | 2.00% | 2.50% | 12.70% | 13.40% | 11.30% | 11.80% | 12.80% | 12.50% | 5.70% |
| DCASE2016 Task2 | LLM-C | 0.00% | 0.00% | 21.43% | 35.71% | 7.14% | 14.29% | 21.43% | 21.43% | 7.14% |
| LibriCount | NAR | 30.40% | 42.10% | 35.90% | 42.50% | 3.50% | 0.40% | 22.00% | 23.20% | 26.20% |
@@
| Vocal Imitations | LLM-C | 15.17% | 5.50% | 1.17% | 1.83% | 18.67% | 11.17% | 17.50% | 3.83% | 5.83% |
| Task | Metric | Whisper- | LTU-AS | SALMONN | SALMONN | Qwen- | Qwen2- | WavLLM | MU- | GAMA |
| LLaMA | 7B | 13B | Audio | Audio | LLaMA | |||||
| Giantsteps Key | LLM-C | 3.00% | 0.50% | 1.50% | 1.00% | 3.00% | 17.00% | 2.00% | 0.50% | 0.00% |
| MTG Top50 | LLM-C | 0.00% | 0.10% | 0.00% | 0.00% | 1.40% | 0.20% | 0.00% | 0.60% | 0.10% |
| MagnaTagATune | LLM-C | 5.50% | 1.00% | 3.50% | 9.00% | 4.50% | 1.50% | 1.00% | 3.50% | 7.50% |
| MTG-Genre | LLM-C | 0.60% | 2.10% | 0.00% | 0.00% | 3.70% | 0.80% | 0.80% | 0.80% | 0.30% |
| GTZAN | LLM-C | 28.50% | 8.00% | 32.00% | 61.50% | 63.50% | 68.50% | 22.50% | 17.00% | 5.00% |
| MTGMoodTheme | LLM-C | 0.10% | 1.50% | 0.00% | 0.00% | 1.30% | 0.50% | 0.10% | 0.00% | 0.10% |
| GTZAN-Rhythm | NAR | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% |
@@
| VocalSet | LLM-C | 8.00% | 1.00% | 11.00% | 6.50% | 6.50% | 15.50% | 8.00% | 8.50% | 2.50% |
| MTG Instrument | LLM-C | 1.00% | 1.00% | 0.90% | 0.00% | 9.70% | 3.30% | 1.10% | 1.00% | 0.60% |
@@ Appendix E LLM-based evaluation
Here, we list the prompt templates used in the evaluation. Figure 6 illustrates the prompt for classification task evaluation, while Figure 7 shows the prompt for regression task evaluation. In addition to GPT-4o, we also employed open-source LLMs to ensure better reproducibility. Tables 11 and 12 present the performance evaluations of the proposed prompts on classification and regression tasks, respectively, comparing their correlation with a human-labeled evaluation set. We found that GPT-4o and LLaMA-3.1-70B-Instruct achieved very high accuracy in both classification and regression tasks. This demonstrates that our evaluation using GPT-4o is reliable and that researchers can leverage LLaMA-3.1-70B-Instruct to conduct fully reproducible evaluations.
Classification Post-processing Prompt:
Please strictly follow the guidelines below:
- Briefly explain the reasons for your judgment.
- Answer with the format ”Result: YES or NO” at the end.
- The questions would be single-choice or multi-choice:
Input Format:
Question: {instruction}
Ground Truth Answer: {label}
Model’s Response: {response}
@@ Appendix G Audio duration distribution
Figure 11 shows the distribution of all audio files in Dynamic-SUPERB Phase-2. Only about 7% of the audio files have a duration longer than 30 seconds, which reflects an inherent limitation of models that incorporate Whisper in their framework.

### v2

@@ Appendix D HEAR & MARBLE Evaluation Results
Tables 7 and 8 show the evaluation results for the HEAR and MARBLE tasks, respectively, as collected in Dynamic-SUPERB Phase-2. Importantly, we replaced the datasets for some tasks to address licensing issues and reduced their size to enable more efficient inference. Furthermore, we removed certain tasks because they overlap with existing SUPERB tasks. Consequently, these results are not directly comparable to those evaluated on the original HEAR and MARBLE.
| Task | Metric | Whisper- | LTU-AS | SALMONN | SALMONN | Qwen- | Qwen2- | WavLLM | MU- | GAMA-IT |
| LLaMA | 7B | 13B | Audio-Chat | Audio-7B-Inst. | LLaMA | |||||
| Beehive States Classification - Beehive States | LLM-C | 43.00% | 35.00% | 54.00% | 42.50% | 19.00% | 42.50% | 37.00% | 37.50% | 19.00% |
| Emotion Recognition - CREMAD | LLM-C | 6.70% | 9.30% | 18.80% | 18.10% | 62.50% | 61.10% | 25.60% | 12.40% | 13.20% |
| Environmental Sound Classification - ESC-50 | LLM-C | 4.80% | 25.30% | 61.00% | 73.40% | 68.80% | 43.90% | 9.80% | 0.40% | 35.30% |
| Language Identification - VoxLingua107 Top10 | LLM-C | 92.18% | 1.95% | 6.28% | 11.01% | 18.00% | 36.11% | 1.85% | 0.00% | 0.10% |
| Music Genre Classification - ISMIR04 | LLM-C | 15.58% | 7.04% | 31.16% | 44.22% | 33.67% | 37.69% | 16.58% | 27.64% | 2.51% |
| Music Speech Classification - MAESTRO-LibriSpeech | LLM-C | 72.50% | 77.50% | 91.50% | 99.50% | 98.00% | 89.50% | 59.50% | 51.50% | 88.00% |
| Music Transcription - MAESTRO | NAR | 92.43% | 100.00% | 99.46% | 100.00% | 96.76% | 96.76% | 96.76% | 100.00% | 100.00% |
@@
| Percussion Instruments Classification - Beijing Opera Percussion Instrument | LLM-C | 3.39% | 4.24% | 10.17% | 20.76% | 16.10% | 27.12% | 24.15% | 22.03% | 11.44% |
| Percussion Instruments Stroke Classification - Mridangam Stroke | LLM-C | 0.10% | 3.80% | 7.60% | 4.90% | 11.50% | 10.80% | 13.70% | 10.30% | 0.10% |
| Percussion Instruments Tonic Classification - Mridangam Stroke | LLM-C | 2.00% | 2.50% | 12.70% | 13.40% | 11.30% | 11.80% | 12.80% | 12.50% | 5.70% |
| Sound Event Detection - DCASE2016 Task2 | LLM-C | 0.00% | 0.00% | 21.43% | 35.71% | 7.14% | 14.29% | 21.43% | 21.43% | 7.14% |
| Speaker Count Identification - LibriCount | NAR | 30.40% | 42.10% | 35.90% | 42.50% | 3.50% | 0.40% | 22.00% | 23.20% | 26.20% |
@@
| Vocal Imitation Classification - Vocal Imitations | LLM-C | 15.17% | 5.50% | 1.17% | 1.83% | 18.67% | 11.17% | 17.50% | 3.83% | 5.83% |
| Task | Metric | Whisper- | LTU-AS | SALMONN | SALMONN | Qwen- | Qwen2- | WavLLM | MU- | GAMA-IT |
| LLaMA | 7B | 13B | Audio-Chat | Audio-7B-Inst. | LLaMA | |||||
| Key Detection - Giantsteps (Key) | LLM-C | 3.00% | 0.50% | 1.50% | 1.00% | 3.00% | 17.00% | 2.00% | 0.50% | 0.00% |
| Music Tagging - MTG-Jamendo | LLM-C | 0.10% | 0.30% | 0.00% | 0.00% | 2.90% | 0.30% | 0.10% | 0.50% | 0.10% |
| Music Tagging - MagnaTagATune | LLM-C | 5.50% | 1.00% | 3.50% | 9.00% | 4.50% | 1.50% | 1.00% | 3.50% | 7.50% |
| Genre Classification - MTG-Jamendo | LLM-C | 0.00% | 1.60% | 0.10% | 0.20% | 3.20% | 0.10% | 0.20% | 0.20% | 0.00% |
| Emotion Detection - MTG-Jamendo | LLM-C | 0.00% | 1.50% | 0.00% | 0.00% | 2.70% | 0.80% | 0.80% | 0.80% | 0.10% |
| Beat Tracking - ASAP | NAR | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% |
@@
| Vocal Technique Detection - VocalSet | LLM-C | 8.00% | 1.00% | 11.00% | 6.50% | 6.50% | 15.50% | 8.00% | 8.50% | 2.50% |
| Instrument Classification - MTG-Jamendo | LLM-C | 0.40% | 2.30% | 0.00% | 0.00% | 6.40% | 1.40% | 0.20% | 0.60% | 0.20% |
@@ Appendix E LLM-based evaluation
Here, we list the prompt templates used in the evaluation. Figure 6 illustrates the prompt for classification task evaluation, while Figure 7 shows the prompt for regression task evaluation. In addition to GPT-4o, we also employed open-source LLMs to ensure better reproducibility. Tables 9 and 10 present the performance evaluations of the proposed prompts on classification and regression tasks, respectively, comparing their correlation with a human-labeled evaluation set. We found that GPT-4o and LLaMA-3.1-70B-Instruct achieved very high accuracy in both classification and regression tasks. This demonstrates that our evaluation using GPT-4o is reliable and that researchers can leverage LLaMA-3.1-70B-Instruct to conduct fully reproducible evaluations.
@@
Classification Post-processing Prompt:
Please strictly follow the guidelines below:
- Briefly explain the reasons for your judgment.
- Answer with the format ”Result: YES or NO” at the end.
- The questions would be single-choice or multi-choice:
Input Format:
Question: {instruction}
Ground Truth Answer: {label}
Model’s Response: {response}
@@ Appendix G Audio duration distribution
Figure 11 shows the distribution of all audio files in Dynamic-SUPERB Phase-2. Only about 6.5% of the audio files have a duration longer than 30 seconds, which reflects an inherent limitation of models that incorporate Whisper in their framework.